# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ V.2 QBCD in the $3$-regular Max Cut problem
As demonstrated above, the QBCD Hamiltonian achieves finite ground-state fidelities at exponentially shorter driving times compared to the first- and second-order local CD expansions, within the minimal spin-glass bottleneck model. This is evidenced by the polynomial closing of the minimum energy gap and by the finite fidelities in the exponentially long plateaus, with vanishing occupation of the ground state with the local CD expansion. Moreover, the QBCD approach proves remarkably robust: its performance remains stable even under significant deviations from the exact critical point. It also retains its effectiveness when constructed through a purely numerical procedure, which requires computational resources that are subleading compared to those needed for simulating the full time evolution. This numerically accessible construction not only improves practicality but also brings experimental realization within closer reach. Furthermore, the QBCD protocol maintains a bounded energy scale, further underscoring its advantage compared to the linearly diverging energy cost of the local expansions.
To this end, we investigate the performance of the local CD expansion and the QBCD method on a realistic NP-hard problem Hamiltonian, the $3$-regular Max Cut [114, 115]. The $3$-regular Max Cut problem describes the optimization task of dividing a $3$-regular random graph into two subgraphs such that the number of connecting edges is maximal. The solution for a given graph is equivalent to finding the ground state of the following Hamiltonian,
$H=\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2},$ (40)
where $c$ denotes the edge connecting the two spins representing the vertices of the graph. Edge configurations are distributed uniformly such that all spins appear in exactly three interaction terms without repetition, $c_{1}\neq c_{2}$, implying a total $M=3L/2$ number of quadratic terms. In Ref. [114], it was shown that the ground state energy gap of the Max Cut exhibits an exponential decay even for intermediate system sizes, $L=8,\dots,18$ for problem instances with two-fold ground state degeneracy. Correspondingly, we also restrict the exposition to these instances and consider the time-evolution initiated from the fully paramagnetic ground state governed by the Hamiltonian
@@
where the time-dependence of the linear protocol was omitted for brevity, $\lambda(t)=t/T,\quad t\in[0,T]$.
First, we probe the efficiency of the local variational expansion. Similar to Sec. III.2, we consider two and three-body terms in which all interactions are restricted to spins connected by the graph edges. Allowing also for all local couplings to be optimized, the first and second order CD corrections read
@@
In the second-order expansion, $H^{(2)}_{1}$, three-body terms run over connected spins in the interaction graph of the Max Cut problem, indicated by the constraint $\delta_{c_{2},c^{\prime}_{1}}$ and the adapted order of indices, $c_{1}<c_{2}=c^{\prime}_{1}<c^{\prime}_{2}$. The corresponding minimization equation with the first order expansion is given by (see App. F)
$\left[32(1-\lambda)^{2}+12\lambda^{2}\right]\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]=-2(\lambda-2).$ (44)
The local parameters for the second order term (for the detailed derivation see App. F) are obtained by minimizing the action, Eq. (26), with respect to both $\alpha_{c}$ and $\beta_{c,c^{\prime}}$,
$\left[32\lambda^{2}+16(1-\lambda)^{2}\right]\alpha_{c}+(1-\lambda)^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left(\delta_{c_{2},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+\delta_{c_{1},c^{\prime}_{2}}\right)$ (45)
$+8(1-2\lambda)\sum_{c^{\prime}}(\beta_{c,c^{\prime}}\delta_{c^{\prime}_{1},c_{2}}+\beta_{c^{\prime},c}\delta_{c^{\prime}_{2},c_{1}})=-2(2-3\lambda),$
$16((1-\lambda)^{2}+\lambda^{2})\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}+\lambda^{2}\sum_{c^{\prime\prime}}\beta_{c^{\prime\prime},c^{\prime}}\left[(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c^{\prime\prime}_{2},c_{2}}+(1-\delta_{c^{\prime\prime}_{2},c_{2}})\delta_{c^{\prime\prime}_{1},c_{1}}\right]=2(2\lambda-1)\delta_{c_{2},c^{\prime}_{1}}(\alpha_{c}+\alpha_{c^{\prime}}).$ (46)
Similar to the case of the minimal spin glass model, the low-order local CD expansion provides a negligible improvement in shortening the adiabatic time scale. Besides the ground state gap, we investigate the ground state fidelity, which is a more suitable measure of adiabaticity for the achievable system sizes, $L=6,\dots,18$. As shown in Fig. 12, the ground state gap increases only by a constant factor, even for short driving times, $T=2$. This increase is slightly more pronounced in the second-order expansion compared to the slow exponential decay observed in the bare process without CD. Additionally, our findings also reveal that the relative difference from the gap without CD follows a universal curve decaying approximately as $(\Delta^{(1,2)}_{\mathrm{min,CD}}-\Delta_{\mathrm{min}})/\Delta_{\mathrm{min}}\sim T^{-2}$, as shown in Fig. 13. Notably, this result further demonstrates the robustness of the predictions drawn from the minimal spin glass model, Fig. 6. Regarding the final ground state fidelities, even weaker efficiency is observed for the first two orders of the local expansion. While for long driving times, minimal improvement is observed, for faster processes, the final fidelity becomes even smaller than in the bare process. These observations are displayed in Fig. 14.
Next, we test the performance of the QBCD Hamiltonian, constructed by the ground state gap, $\Delta$, and the many-body ground ($\lvert\mathrm{GS}\rangle$) and first excited states ( $\lvert\mathrm{ES}\rangle$) near the critical point
@@
$-i\frac{\langle\mathrm{GS}\lvert\partial_{\lambda}H\rvert\mathrm{ES}\rangle}{\Delta}\left(\lvert\mathrm{ES}\rangle\langle\mathrm{GS}\rvert-\lvert\mathrm{GS}\rangle\langle\mathrm{ES}\rvert\right).$
The approximate knowledge of these quantities is obtained via the same purely numerical procedure employed for the minimal spin glass model, based on a search over only $40$ parameter values. While in the minimal model this choice had limited computational relevance, for the realistic spin glass problem, it clearly requires only subleading numerical costs compared to the subsequent QBCD-assisted simulation.
To further improve its feasibility for near-term experimental realizations, we sparsify the QBCD Hamiltonian to the density of the quadratic local expansion by keeping only the $L(L+1)/2$ largest absolute valued matrix elements. Although the exact QBCD construction involves only two states, selected out of the $\sim 4^{L}$ possible energy level pairs of the full CD driving, the sparsification further reduces its non-locality exponentially, bringing its experimental realization within reach. Moreover, the computational cost of constructing this sparsified QBCD Hamiltonian grows only polynomially with the system size, making it accessible for classical large-scale algorithms as well [116, 117].
To this end, the performance of the sparsified QBCD is also tested via the ground state gap and the final fidelity. As shown in Fig. 12, the ground state gap is exponentially amplified, exhibiting an approximately polynomial decay even at intermediate driving times, $T=5,20$, whereas it only approaches the behavior without CD driving close to the fully adiabatic timescales. The fidelity exhibits similar features, remaining finite with the sparsified QBCD even at short driving times, $T=1,5$, while the fidelity achieved using the local CD approach becomes exponentially small under the same conditions. In the latter case, the fidelity only becomes finite close to adiabaticity, but with negligible improvements compared to the bare process. Whereas, in this regime, the sparsified QBCD performs more efficiently, achieving approximately unit fidelity.
@@
Figure 13: Relative difference between ground-state gaps without CD and with the local CD approaches. The improvement achieved by first- (solid lines, filled dots) and second-order (dashed lines, empty dots) local expansions decays approximately quadratically with system size. Notably, the amplification of the gap vanishes at driving times just preceding the onset of finite fidelity.
@@
Finally, we compare the performance of the sparsified QBCD and the local expansion with the COLD approach [77]. Within this scheme, one introduces an additional local operator, $\mathcal{O}_{\mathrm{opt}}$, to the problem Hamiltonian with optimizable additional parameters,
@@
$-\lambda\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j=1}^{L}\,\sigma^{x}_{j}+f(t,\boldsymbol{\gamma})\mathcal{O}_{\mathrm{opt}}.$
For computational convenience and following the strategy of Ref. [77], we choose the optimized operator as the sum of longitudinal spin operators, $\mathcal{O}_{\mathrm{opt}}=\sum_{j}\sigma^{z}_{j}$. The time- and parameter-dependent coefficient is represented by its Fourier decomposition,
$f(t,\boldsymbol{\gamma})=\sum_{k=1}^{N_{k}}\gamma_{k}\sin\left(\pi\frac{t}{T}\right),$ (49)
with $N_{k}$ denoting the number of modes. This form ensures that the CD term vanishes at the initial and final times, i.e., $f(t=0,T,\boldsymbol{\gamma})=0$, thereby preserving the original Hamiltonian at the endpoints. The Fourier coefficients, $\gamma_{k}$ are numerically optimized within the local CD expansion framework, where the final fidelity of the evolved state serves as the objective function. Since the presence of the longitudinal field allows for linear Pauli operators in the lowest order CD expansion, we consider the second-order local CD approach
$H^{(2)}_{\mathrm{COLD}}=\sum_{j}\epsilon_{j}\sigma^{y}_{j}+\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}).$ (50)
@@
The minimization principle is written in terms of this second-order local CD term, together with the optimizable operator,
[…]
$G(H^{(2)}_{\mathrm{COLD}})=$ $\partial_{t}H(t)+\partial_{t}f(t,\boldsymbol{\gamma})\mathcal{O}_{\mathrm{opt}}$
$-i[H(t),H^{(2)}_{\mathrm{COLD}}]-if(t,\boldsymbol{\gamma})[\mathcal{O}_{\mathrm{opt}},H^{(2)}_{\mathrm{COLD}}].$
@@
The corresponding trace minimization and the resulting system of equations for the sets of $\alpha_{c}\left(\lambda,f(t,\boldsymbol{\gamma})\right)$ and $\epsilon_{j}\left(\lambda,f(t,\boldsymbol{\gamma})\right)$ are detailed in App. G. To perform the numerical optimization, we employed MATLAB R2023’s built-in genetic algorithm to search for the optimal set of Fourier coefficients $\gamma_{k}$ that maximize the final fidelity. The optimization was repeated $100$ times to ensure the convergence of the results. As a first clear sign of the inferiority of the COLD approach compared to the QBCD, the procedure proved to be computationally demanding, restricting our analysis to systems with up to $L=10$ spins—only almost half the size studied in the QBCD framework. Due to similar numerical restrictions, the number of Fourier modes was also limited to $N_{k}=4$. Despite these constraints, our setting already surpasses that of Ref. [77], by using twice the number of spins and one extra Fourier mode, allowing for a more systematic exploration of its performance in spin glasses. Crucially, the fact that such modest extensions already impose severe computational bottlenecks underscores a central message: the practical implementability and efficiency of the QBCD method is fundamentally superior to what the COLD approach can offer. They also highlight the significant advantage of the QBCD approach—not only in terms of algorithmic scalability but also in terms of its greater feasibility for experimental realization.
To this end, we benchmarked the performance of the second-order COLD method against both the sparsified QBCD protocol and the second-order local CD expansion, Eq. (43), using the final ground state fidelity as the primary measure for $L=6,8,10$ spins. Given the small system size, averaging over $10$ random graph instances was sufficient to obtain well-converged results for the average fidelity. As shown in Fig. 15, the COLD approach clearly outperforms the local expansion, particularly at intermediate annealing times, reflecting the benefits of its variationally optimized structure. Although the fidelity decreased rapidly in faster processes, it converged to the local expansion and the bare process as the adiabatic limit is reached.
Notably, the advantage of QBCD becomes most prominent modest driving regime, $T\lesssim 15$, where it continues to maintain fidelities close to unity. In contrast, both local and the COLD approaches begin to deteriorate, highlighting the onset of nonperturbative spin-glass effects that resist purely local counterdiabatic corrections. These results compellingly demonstrate that even in small systems, the spin-glass bottleneck can severely limit the effectiveness of local methods, regardless of how the driving protocol is optimized.
Note that the relatively small system size allows optimization effects to be more visible. In this regime, the COLD protocol—despite being limited to two-body operators—can slightly outperform the second-order local expansion with three-body terms. However, this advantage starts to fade away as the system size increases. The emergence of stronger spin-glass behavior and a growing number of diabatic transitions lowers the performance of COLD toward that of the second-order local expansion. Meanwhile, the superiority of QBCD becomes increasingly pronounced, as it continues to capture essential global features—through its exponentially suppressed non-locality—that remain out of reach for purely local approaches.
@@
Figure 15: Fidelity as a function of the driving time for the sparsified QBCD, COLD approach, second order local expansion, and for the bare process, for various system sizes. (a) Even for $L=6$ QBCD provides a sizable improvement at intermediate times. (b) The advantage of QBCD becomes more apparent also at longer driving times for $L=8$. (c) For $L=10$ even close to adiabaticity, the QBCD clearly outperforms the COLD, while the latter gets closer to the second-order expansion and the bare process results.
V.3 Demonstration on the $3$-regular $3$-XORSAT

### v3

@@ V.2 QBCD in the $3$-regular Max Cut problem
Motivated by the efficiency of the QBCD approach in the minimal spin-glass bottleneck, we now consider a more realistic NP-hard problem, the $3$-regular Max Cut [115, 116], to assess QBCD in a practical setting. The $3$-regular Max Cut problem describes the optimization task of dividing a $3$-regular random graph into two subgraphs such that the number of connecting edges is maximal. The solution for a given graph is equivalent to finding the ground state of the following Hamiltonian,
$H_{P}=\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2},$ (30)
where $c$ denotes the edge connecting the two spins that represent the vertices of the graph. Edge configurations are distributed uniformly so that all spins appear in exactly three interaction terms without repetition, $c_{1}\neq c_{2}$, implying a total $M=3L/2$ number of quadratic terms. In Ref. [115], it was shown that the ground state energy gap of Max Cut exhibits an exponential decay even for intermediate system sizes, $L=8,\dots,18$ for problem instances with two-fold ground state degeneracy. Correspondingly, we also restrict the exposition to these instances and consider the time evolution initiated from the fully paramagnetic ground state governed by the Hamiltonian
@@
where the time dependence of the linear protocol was omitted for brevity, $\lambda(t)=t/T,\quad t\in[0,T]$.
First, we investigate the efficiency of the local variational expansion. Similarly to Sect. III.2, we consider two- and three-body terms in which all interactions are restricted to spins connected by the graph edges. Allowing also for all local couplings to be optimized, the first- and second-order CD corrections read
@@
In the second-order expansion, $H^{(2)}_{1}$, three-body terms run over connected spins in the interaction graph of the Max Cut problem, indicated by the constraint $\delta_{c_{2},c^{\prime}_{1}}$ and the adapted order of indices, $c_{1}<c_{2}=c^{\prime}_{1}<c^{\prime}_{2}$. The corresponding minimization equation for the first- and second-order expansions can be found in App. F.
To assess adiabaticity for system sizes $L=6,\dots,18$, we consider both the ground-state gap and fidelity. As shown in Fig.11, the variational approach can amplify the gap only by a constant factor, even for short driving times, $T=2$ (with adiabaticity completely broken down due to excitations to higher energy modes), which is slightly more pronounced for the second-order expansion compared to the slow decay in the bare process; see App.G. The final ground-state fidelities show even weaker gains: for slow protocols, there is minimal improvement, while for faster processes, fidelities can drop below those of the bare evolution (Fig.12). In contrast to this modest gap amplification and limited fidelity gains observed, QBCD retains its efficiency and robustness in these more realistic spin-glass models, similarly to the minimal spin-glass bottleneck. In particular, the QBCD Hamiltonian is constructed using the ground state gap, $\Delta$, and the many-body ground ($\lvert\mathrm{GS}\rangle$) and first excited states ( $\lvert\mathrm{ES}\rangle$) near the critical point
@@
[…]
The approximate knowledge of these quantities is obtained via the same purely numerical procedure employed for the minimal spin glass model, based on a search over only $40$ parameter values.
To improve its practicality, we sparsify the QBCD Hamiltonian by retaining only the $L(L+1)/2$ largest matrix elements, matching the density of the quadratic local expansion. Although the exact QBCD construction involves only two states, selected out of the $\sim 4^{L}$ possible energy level pairs of the full CD driving, the sparsification further reduces its non-locality exponentially, enabling a practically realizable approximate implementation through controlled truncation. This truncated scheme substantially alleviates the resource demands of current local CD protocols and can facilitate the efficient realization of higher-order CD expansions on quantum devices. Moreover, the computational cost of constructing this sparsified QBCD Hamiltonian grows only polynomially with the system size, making it accessible for classical large-scale algorithms as well [117, 118].
The sparsified QBCD is tested via the ground-state gap and final fidelity. As shown in Fig. 11, the gap is exponentially amplified, decaying only approximately polynomially even at intermediate driving times ($T=5,20$) and approaching the bare behavior near fully adiabatic timescales. Correspondingly, the final fidelity remains finite at short driving times ($T=1,5$), whereas the variational approach yields exponentially small fidelity except near adiabaticity, highlighting that sparsified QBCD simultaneously amplifies the gap and preserves high fidelity even far from the adiabatic regime. Thus, our results demonstrate that sparsified QBCD pinpoints an exponentially small non-local fraction whose inclusion alone suffices to achieve exponential speedups. Although extrapolation to significantly larger system sizes remains challenging, the underlying key microscopic mechanisms indicate that comparable performance may also extend to a broader range of instances.
@@
We also compare the performance of the sparsified QBCD and the local expansion with the COLD approach [77]. Within this scheme, one introduces an additional local operator, $\mathcal{O}_{\mathrm{opt}}(\boldsymbol{\gamma},t)$, to the problem Hamiltonian with optimizable additional parameters,
@@
$-\lambda\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j=1}^{L}\,\sigma^{x}_{j}+\mathcal{O}_{\mathrm{opt}}(t,\boldsymbol{\gamma}).$
Following the strategy of Ref. [119], we choose $\mathcal{O}_{\mathrm{opt}}(t,\boldsymbol{\gamma})$ as the second-order nested commutator $\mathcal{O}_{\mathrm{opt}}(t,\boldsymbol{\gamma})=\gamma^{(1)}\sin(2\pi\lambda)\left[H_{0}\left[H_{1},H_{0}\right]\right]+\gamma^{(2)}\sin(2\pi\lambda)\left[H_{1}\left[H_{1},H_{0}\right]\right]$, providing a more efficient annealing path by extending the total Hamiltonian with the even nested commutators. The time- and parameter-dependent coefficients, $\gamma^{(1,2)}$, are subjected to further optimization within the local CD expansion framework, where the final fidelity of the evolved state serves as the objective function. Since the double commutator structure will contain terms like $\left[H_{0}\left[H_{1},H_{0}\right]\right]\sim ZXZ+X$ and $\left[H_{1}\left[H_{1},H_{0}\right]\right]\sim YY+ZZ$, the general form of the local CD expansion up to cubic terms matches that of Eq. (33)
$H^{(2)}_{\mathrm{COLD}}$= $\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}})$
$+\sum_{c,c^{\prime}=1}^{M}\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}).$
@@
The minimization principle is written in terms of this local CD term, together with the optimizable operator,
$G(H^{(2)}_{\mathrm{COLD}})$= $\partial_{t}H(t)+\partial_{t}\mathcal{O}_{\mathrm{opt}}(t,\boldsymbol{\gamma})$
$-i[H(t),H^{(2)}_{\mathrm{COLD}}]-i[\mathcal{O}_{\mathrm{opt}}(t,\boldsymbol{\gamma}),H^{(2)}_{\mathrm{COLD}}$].
@@
The corresponding trace minimization and the resulting system of equations for the sets of $\alpha_{c}\left(\lambda,t,\boldsymbol{\gamma}\right)$ and $\beta_{c,c^{\prime}}\left(\lambda,t,\boldsymbol{\gamma}\right)$ are detailed in App. H. To perform the numerical optimization, we employed MATLAB R2023’s built-in genetic algorithm to search for the optimal set of Fourier coefficients $\gamma^{(1,2)}$ that maximizes the final fidelity. The optimization was repeated $100$ times to ensure the convergence of the results. As a first clear sign of the inferiority of the COLD approach compared to the QBCD, the procedure proved to be computationally demanding, restricting our analysis to systems with up to $L=10$ spins—only almost half the size studied in the QBCD framework, which already exceeded the QBCD running time by several orders of magnitude for the largest system sizes. Despite these constraints, our setting already surpasses that of Ref. [77, 119], by using twice the number of spins, allowing for a more systematic exploration of its performance in spin glasses. These results consistently show that even modest extensions of local CD methods rapidly incur severe computational bottlenecks while providing only marginal benefits in maintaining adiabaticity near the critical point, reinforcing the conclusion that QBCD offers a far more practical and scalable route for larger systems.
Furthermore, we benchmarked the COLD method against the sparsified QBCD protocol and the second-order local CD expansion, Eq. (33), using the final ground-state fidelity for $L=6,10$ spins averaged over 10 random graph instances. As shown in Fig. 13, COLD benefits from its variational optimization at intermediate and long annealing times. For small systems, COLD can get close to the sparsified QBCD method, but this advantage diminishes with increasing system size due to stronger spin-glass effects and additional diabatic transitions. QBCD, in contrast, maintains near-unit fidelity in the modest driving regime ($T\lesssim 15$) by capturing global features and suppressing non-local corrections, consistently outperforming local CD approaches.
Finally, we compare the energetic requirements of the local CD expansion with those of the sparsified QBCD, by studying the system size dependence of the Hilbert-Schmidt norm, similar to the minimal spin-glass model [Sec. V.1]. As demonstrated in Fig. 16, for the Max Cut it implies the following scalings:
$\lvert\lvert H^{(1)}_{1}\rvert\rvert^{2}\sim L\qquad\text{and}\qquad\lvert\lvert H^{(2)}_{1}\rvert\rvert^{2}\sim L^{2}.$ (38)
By contrast, the sparsified QBCD displays a much slower growth, following approximately a linear slope for $L\lesssim 20$, owing to the exponentially reduced nonlocality and the finite deviations from the exact critical point (see Fig. 16 as well). While the divergence of the energy denominator is significantly weakened by the finite deviations from the critical point in the approximate numerical construction, the numerator itself becomes exponentially small. This suppression stems from the sparse structure of $\partial_{\lambda}H$, the eigenstates, and their overall normalization factor scaling as $\sim 2^{-L}$. Although the scaling is only approximately linear, the underlying microscopic mechanisms indicate that such moderate growth is a comprehensible scenario for larger systems. Taken together, they demonstrate that sparsified QBCD not only avoids the critical-point divergence but also achieves a more favorable energetic scaling while still retaining its efficiency in overcoming spin-glass bottlenecks.This further demonstrates how QBCD provides a new avenue for approximate CD, facilitating implementations that place substantially lower demands on external resources through systematic truncation.
@@
Figure 13: Fidelity as a function of the driving time for the sparsified QBCD, COLD approach, second-order local expansion, and for the bare process, for various system sizes. (a) Even for $L=6$ QBCD provides a sizable improvement at intermediate times. (b) For $L=10$ even close to adiabaticity, the QBCD clearly outperforms the COLD, while the latter gets closer to the second-order expansion and the bare process results.
V.3 Demonstration on the $3$-regular 33-XORSAT